OpenStudy (rayman5766):

A bag contains 6 nickels and 14 pennies. What is the ratio, written in lowest terms, of nickels to the total number of coins in the bag?

OpenStudy (anonymous):

6:14 u know how to reduce?

OpenStudy (anonymous):

just reduce that \[\frac{ 6 }{ 14 }\]

OpenStudy (anonymous):

first find no. of coins.

OpenStudy (rayman5766):

@TheRaggedyDoctor my dump computer restarted and now I have to do it over again my god I am so angry

OpenStudy (rayman5766):

I cant rember I am so angry right now was it 6/3

OpenStudy (anonymous):

then reduce 6/?

OpenStudy (theraggedydoctor):

http://openstudy.com/users/rayman5766 You can just go to that link, open some of the questions you asked, and find the answers I gave you :)

OpenStudy (anonymous):

14\[\div6\]

OpenStudy (anonymous):

14/6=?

OpenStudy (rayman5766):

ok thanks

OpenStudy (anonymous):

no total no. of coins =6+14=20 then 6/20 reduce it to lower terms.

OpenStudy (rayman5766):

3/10

OpenStudy (anonymous):

correct.